# Probabilistic modeling of Cherenkov emission from particle showers

### v2

@@ Abstract
Subatomic particles can interact with target nuclei in matter or decay in flight, and an individual high-energy particle can induce a particle shower composed of numerous, lower-energy secondaries. These particle showers broadly exhibit universality across diverse media, including air, water, ice, and other materials, with their development governed by the Standard Model. Full Monte Carlo simulation of particle showers, where each secondary is individually tracked and propagated, can be a computational challenge to perform at scale. Experiments thus resort to parametrized approximations when efficient simulation becomes necessary. Here, we construct distributions of parameters capable of describing the Cherenkov light yield from particle showers in ice, and extensible to other, similar media. Sampling from the distributions allows for a much improved description of event-to-event fluctuations, in amplitude and shape, along the shower axis. Including these effects is essential for a more accurate simulation of signal and background events in current and next-generation neutrino telescopes.
@@ 2.2 Characteristics of FLUKA results
The rationale for such a transformation will be discussed in section˜3.1. The same line and color style is used as in the left panel, corresponding to different primary particles. Values of $\varrho(a^{\prime},b^{\prime})$ lie close to one for all primaries, indicating strong correlation. In contrast, $\varrho(\hat{\ell}_{\text{tot}},\{a^{\prime},b^{\prime}\})$ for EM showers lie near zero, indicating a lack of correlation, while those for hadron primaries exhibit some anticorrelation as energy increases. For hadrons below roughly $10\text{\,}\mathrm{GeV}$, the primary particle loses energy via ionization and excitation of electrons. Decays also occur with increasing probability for the mesons and heavier baryons. The resulting Cherenkov yield from these dominant, stochastic processes are difficult to describe using the parametrization in section˜3. Fully MC-based approaches are recommended instead, especially as the computational needs are also lower.
@@ 4 Application to high-energy neutrino interactions
A more comprehensive visualization is presented in figure˜10 in appendix˜D. There, $\mathrm{8}\mathrm{0}$ $\text{d}\hat{\ell}/\text{d}x$ outcomes are shown, with the first two corresponding to that of figure˜9. While NC interactions (blue panels) tend to present more structure at lower amplitudes, CC interactions exhibit variation in terms of the overall extension and, in a few cases, additional structure is visible.
The example discussed in this section is focused on electron-neutrino DIS, but the shower parametrizations are applicable to other processes as well. The stochastic energy loss of high-energy muons typically proceeds via Bremsstrahlung, pair production or photonuclear interactions, all of which result in primaries that initiate particle showers. Decays of tau leptons or $W$ bosons can also produce such primaries. More generally, for any high-energy simulation interested in the Cherenkov light yield of particle showers, physics processes can be tracked using MC up to a certain threshold. Beyond that, a switch to using the parametrizations detailed in section˜3 can yield a substantial speed-up in the simulation.
@@ 5 Discussion
Several simplifications were assumed in the development of this model, which future work can improve upon. The most obvious of these is that only a one-dimensional $\text{d}\hat{\ell}/\text{d}x$ is modeled, and a fully three-dimensional model likely would require new methods. Even in one dimension, the gamma distribution is an approximation and does not capture multiple peaks, which may arise due to a combination of multiple subshower components. A couple examples of outliers are shown in the left panel of figure˜7. The factorization of shape parameters $(a^{\prime},b^{\prime})$ from the amplitude $\hat{\ell}_{\text{tot}}$ also neglects any shape-amplitude correlations. This may be a reasonable assumption for most of the energy ranges studied here, but breaks down in cases such as decay. Examples of $\hat{\ell}_{\text{tot}}$ distributions at lower energies, where decay begins to dominate, are shown in the right panel of figure˜7. Finally, muons from hadronic decays were not included as discussed in section˜2.1. Typically, these muons do not carry away large fractions of the primary energy, and can be handled separately, but could be included in a more complete model.
@@ Appendix D Additional results for $100\text{\,}\mathrm{TeV}$ electron neutrino DIS
Figure˜10 shows $\mathrm{8}\mathrm{0}$ $\text{d}\hat{\ell}/\text{d}x$ profiles of $100\text{\,}\mathrm{TeV}$ electron-neutrino interactions, as discussed in section˜4. The first two correspond to that shown in figure˜9. Each neutrino interacts off a fixed proton target, with hadronization and decays of particles up to $c\tau_{0}<$0.5\text{\,}\mathrm{mm}$\$ simulated with PYTHIA8 Sjöstrand et al. (2015). The resulting final state $e^{-}$, $\gamma$ and hadrons are passed to the model of section˜3, which is used to sample $\text{d}\hat{\ell}/\text{d}x$ according to their energies. In figure˜10, panels highlighted in blue correspond to NC interactions, otherwise CC.

### v3

@@ Abstract
Subatomic particles can interact with target nuclei in matter or decay in flight, and an individual high-energy particle can induce a particle shower composed of numerous, lower-energy secondaries. These particle showers broadly exhibit universality across diverse media, including air, water, ice, and other materials, with their development governed by the Standard Model. Full Monte Carlo simulation of particle showers, where each secondary is individually tracked and propagated, can be a computational challenge to perform at scale. Experiments thus resort to parametrized approximations when efficient simulation becomes necessary. Here, we construct distributions of parameters capable of describing the Cherenkov light yield from particle showers in ice or water. Sampling from the distributions allows for a much improved description of event-to-event fluctuations, in amplitude and shape, along the shower axis. Including these effects is essential for a more accurate simulation of signal and background events in current and next-generation neutrino telescopes.
@@ 2.2 Characteristics of FLUKA results
The rationale for such a transformation will be discussed in section˜3.1. The same line and color style is used as in the left panel, corresponding to different primary particles. Values of $\varrho(a^{\prime},b^{\prime})$ lie close to one for all primaries, indicating strong correlation. In contrast, $\varrho(\hat{\ell}_{\text{tot}},\{a^{\prime},b^{\prime}\})$ for EM showers lie near zero, indicating a lack of correlation, while those for hadron primaries exhibit some anticorrelation as energy increases. For hadrons below roughly $10\text{\,}\mathrm{GeV}$, the primary particle loses energy via ionization and excitation of electrons. Decays also occur with increasing probability for the mesons and heavier baryons. The resulting Cherenkov yield from these dominant, stochastic processes are difficult to describe using the parametrization in section˜3, and to be safe a minimum energy of $100\text{\,}\mathrm{GeV}$ ($10\text{\,}\mathrm{GeV}$) is recommended for hadrons (EM), though the support extends down from those values by an order of magnitude. At these energies, fully MC-based approaches are recommended instead, especially as the computational needs are also lower.
@@ 4 Application to high-energy neutrino interactions
A more comprehensive visualization is presented in figure˜11 in appendix˜D. There, $\mathrm{8}\mathrm{0}$ $\text{d}\hat{\ell}/\text{d}x$ outcomes are shown, with the first two corresponding to that of figure˜9. While NC interactions (blue panels) tend to present more structure at lower amplitudes, CC interactions exhibit variation in terms of the overall extension and, in a few cases, additional structure is visible.
Figure 10: Same as figure˜9 but for $100\text{\,}\mathrm{TeV}$ tau-neutrino CCDIS in ice and x-axis extended up to $30\text{\,}\mathrm{m}$. In the left (right) panel, the outgoing $\tau^{-}$ decays to two $\pi^{0}$ ($\pi^{-}$) and a $\pi^{-}$ ($\pi^{+}$), in addition to a $\nu_{\tau}$. The hadrons, or their immediate decay products, are input to the $\text{d}\hat{\ell}/\text{d}x$ model. The first peak in the left panel is due to hadrons produced in the initial DIS interaction. The $\tau^{-}$ decay lengths are indicated in the panels.
The examples discussed so far focuses on electron-neutrino DIS, but the shower parametrizations are applicable to other processes as well. High-energy muons typically lose energy stochastically via Bremsstrahlung, pair production or photonuclear interactions, all of which result in primaries that initiate particle showers. Decays of tau leptons or $W$ bosons can also produce such primaries. To illustrate, figure˜10 shows the $\text{d}\hat{\ell}/\text{d}x$ distribution for two $100\text{\,}\mathrm{TeV}$ tau-neutrino CCDIS interactions, where both the initial interaction and subsequent tau decay are simulated with PYTHIA8. The left (right) panel corresponds to a scenario where the outgoing $\tau^{-}$ travels $1242\text{\,}\mathrm{cm}$ (88 $8\text{\,}\mathrm{cm}$) and then decays to two $\pi^{0}$ ($\pi^{-}$) and a $\pi^{-}$ ($\pi^{+}$), in addition to a $\nu_{\tau}$. The hadrons, or their immediate decay products in the case of $\pi^{0}$, are inputs for the model, from which $\text{d}\hat{\ell}/\text{d}x$ distributions are then sampled. In the left panel, both the initial DIS and tau decay are visible as two distinct peaks in the $\text{d}\hat{\ell}/\text{d}x$ distribution. Note that stochastic energy losses along the tau track are not simulated, but should be a relatively small contribution at energies below $100\text{\,}\mathrm{TeV}$. The simulation of such stochastic processes for taus and muons can be performed using software such as PROPOSAL Koehne et al. (2013).
More generally, for any high-energy simulation interested in the Cherenkov light yield of particle showers, physics processes can be tracked using MC up to a certain threshold. Beyond that, a switch to using the parametrizations detailed in section˜3 can yield a substantial speed-up in the simulation. The inverse problem of reconstruction can also benefit from a more accurate modeling of Cherenkov emission. In particular, elongation or fluctuations of the shower profile can improve directional reconstruction, if the reconstruction algorithm is flexible enough to fit a superposition of point-like showers Hallen (2013); Abbasi and others (2024). Such an approach may also result in more accurate energy reconstructions, by fitting the individual sub-shower energies. Further, machine learning models trained on simulation would benefit from increased simulation fidelity. Using modern tools, it may be possible to distinguish, on a statistical basis, showers which are more EM-dominated, as seen in figure˜11.
@@ 5 Discussion
Several simplifications were assumed in the development of this model, which future work can improve upon. The most obvious of these is that only a one-dimensional $\text{d}\hat{\ell}/\text{d}x$ is modeled, and a fully three-dimensional model likely would require new methods. Even in one dimension, the gamma distribution is an approximation and does not capture multiple peaks, which may arise due to a combination of multiple sub-shower components. A couple examples of outliers are shown in the left panel of figure˜7. The factorization of shape parameters $(a^{\prime},b^{\prime})$ from the amplitude $\hat{\ell}_{\text{tot}}$ also neglects any shape-amplitude correlations. This may be a reasonable assumption for most of the energy ranges studied here, but breaks down in cases such as decay. Examples of $\hat{\ell}_{\text{tot}}$ distributions at lower energies, where decay begins to dominate, are shown in the right panel of figure˜7. Finally, muons from hadronic decays were not included as discussed in section˜2.1. Typically, these muons do not carry away large fractions of the primary energy, and can be handled separately, but could be included in a more complete model.
@@ Appendix D Additional results for $100\text{\,}\mathrm{TeV}$ electron neutrino DIS
Figure˜11 shows $\mathrm{8}\mathrm{0}$ $\text{d}\hat{\ell}/\text{d}x$ profiles of $100\text{\,}\mathrm{TeV}$ electron-neutrino interactions, as discussed in section˜4. The first two correspond to that shown in figure˜9. Each neutrino interacts off a fixed proton target, with hadronization and decays of particles up to $c\tau_{0}<$0.5\text{\,}\mathrm{mm}$\$ simulated with PYTHIA8 Sjöstrand et al. (2015). The resulting final state $e^{-}$, $\gamma$ and hadrons are passed to the model of section˜3, which is used to sample $\text{d}\hat{\ell}/\text{d}x$ according to their energies. In figure˜11, panels highlighted in blue correspond to NC interactions, otherwise CC.